Find the indicated vaules for the function
f(3)=4x-7

Answer by jim_thompson5910(35256)

Assuming you mean f(x)=4x-7


So f(3)=4(3)-7=12-7=5


So f(3)=5


If you need more help, email me at jim_thompson5910@hotmail.com

Also, feel free to check out my tutoring website

Jim